OpenStudy (anonymous):

[PRECALC] Find a number k such that the given equation has exactly one real solution. x^2+kx+64=0

OpenStudy (anonymous):

takes two steps what is the square root of \(64\) ?

OpenStudy (anonymous):

sqrt(64) = 8

OpenStudy (anonymous):

k good and what is two times eight?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

finished

OpenStudy (anonymous):

\[x^2+16x=64=(x+8)^2\] and \[(x+8)^2=0\] has only one solution , namely \(x=8\)

OpenStudy (anonymous):

That makes sense.